Find the dimension of a rectangular Persian rug whose perimeter is 18 ft. and whose area is 20 ft^2?
what are the length and width?

Answer by stanbon(75887)

Find the dimension of a rectangular Persian rug whose perimeter is 18 ft. and whose area is 20 ft^2?
what are the length and width?
-------------------------------------
Perimeter = 2(L + W)
----------------
L+W = 9
LW = 20
-----
L = 9-W
----
Substitute for "L" and solve for "W":
(9-W)W = 20
-W^2+9W = 20
---
W^2-9W+20 = 0
-----
(W-5)(W-4) = 0
If W = 5, L = 4
If W = 4, L = 5
========================
Cheers,
Stan H.
=============